OpenStudy (anonymous):

Given the following representation of a floating number, what is the equivalent decimal number considering 7bits for exponent and 16bits for mantissa? 0 | 1000111 | 1100110010000000

OpenStudy (anonymous):

i'm guessing the first bit is the sign, do you know what the exponent bias is? i'm going to assume the exponent bias is 64 Exponent: 1000111 = 1 + 1*2 + 1*2^2 + 0*2^3 + 0*2^4 + 0*2^5 + 1*2^6 1000111 = 1 + 2 + 4 + 0 + 0 + 0 + 64 1000111 = 71 and with a bias of 64 the exponent is 71 - 64 = 7 Mantissa: 1100110010000000 = 52352 (i cheated =P but hopefully you can work it out =)) so you should get 52352 * 2^7 = 6701056 hope this helps